D FG Problem 6. A length of wire is cut in half and two length are wrapped together side by side to make a thicker wire. How does the resistance of this new combination compare to the resistance of the origin wire?

Please Help me with my homework problems.
Answer #1

here, let the initial resistance is R

R = p * L/A

L is the length and A is the area of cross section

Now, when the wire is cut in half

R = p * (L/2)/(2A) *

R = 1/4 * p * L/A

hence , the new resistance is 1/4 times the initial resistance